Linda is 6 years younger than Andy. The sum of their ages is 32. What is the equation to this problem if we are trying to find Linda's age?
Answer by lenny460(1073)

Let the age of Andy be n
Therefore the age of Linda is: (n - 6)
Their total ages is: n + (n - 6) = n + n - 6
= 2n - 6 = 32
Add 6 to each side of the equation
2n - 6 + 6 = 32 + 6
Combine like terms:
-6 + 6 = 0
32 + 6 = 38
2n = 38
Divide each side by 2
2n/2 = 38/2
n = 19
n - 6 = 19 - 6
19 - 6 = 13
Andy is 19 years old and
Linda is 13 years old.
19 + 13 = 32